The resistance of aluminium wire of 2 meter length is 7 ohm.What will be the resistance of wire of 8 metre length and of same cross - section area
1
Expert's answer
2012-06-07T10:17:50-0400
R = p · l / A
p2 = p1, A2 = A1, l2 = 4 · l1
So, R2 = 4 · R1 = 4 · 7 = 28
Ohm.